# A general approach to asymptotic elimination of aggregation functions and generalized quantifiers

### v3

@@ 5. Asymptotic equivalence to basic formulas
A commonly considered context is when $D_{n}=\{1,\ldots,n\}$ and $\mathbf{W}_{n}$ is the set of all $\sigma$-structures with domain $D_{n}$. However we want to allow for more generality. For example one may be interested in $\sigma$-structures with domain $D_{n}$ such that all these structures have some common properties. For example, we may retrict some relation symbols in $\sigma$ to be interpreted as relations with degree111 A relation has degree at most $d$ if the Gaifman graph (also called the primal graph) corresponding to the relation has no vertex with degree more than $d$. at most 20 (say), or we could restrict a binary relation symbol $R\in\sigma$ to be interpreted as a tree, a partial order, an equivalence relation, perhaps with some additional properties. We can, but need not, require that, for some $\sigma^{\prime}\subset\sigma$ every structure in $\mathbf{W}_{n}$ has the same reduct to $\sigma^{\prime}$. If such restrictions are imposed on the $\sigma$-structures considered with domain $D_{n}$, then the sequence of cardinalities $|D_{n}|$, $n\in\mathbb{N}^{+}$, may not contain all (sufficiently large) positive integers. Note also that we do not require that $D_{n}\subseteq D_{n+1}$.
With this quite general set-up our goal is to isolate properties (given by Assumption 5.5) of “simpler” 0/1-valued sublogics $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ such that if $L_{0}$ and $L_{1}$ have these properties, then a formula $\varphi(\bar{x})\in PLA^{*}(\sigma)$ is asymptotically equivalent (see Definition 5.2) to an “$L_{0}$-basic formula”, provided that every aggregation function in $\varphi$ is up-continuous (or ct-continuous) with respect to some parameters which are determined only by $(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$ and the subformulas of $\varphi$. By an “$L_{0}$-basic formula” we mean a formula of the form $\bigwedge_{i=1}^{k}(\varphi_{i}(\bar{x})\to c_{i})$ where $\varphi_{i}\in L_{0}$ and $c_{i}\in[0,1]$ for all $i$ and $\lim_{n\to\infty}\mathbb{P}_{n}\big{(}\forall\bar{x}\bigvee_{i=1}^{k}\varphi_{i}(\bar{x})\big{)}=1$. If some $\psi(\bar{x})\in PLA^{*}(\sigma)$ is asymptotically equivalent to an $L_{0}$-basic formula $\varphi(\bar{x})=\bigwedge_{i=1}^{k}(\varphi_{i}(\bar{x})\to c_{i})$, then, for sufficiently large $n$, $\mathcal{A}\in\mathbf{W}_{n}$ and $\bar{a}\in(D_{n})^{|\bar{x}|}$, the value $\mathcal{A}(\psi(\bar{a}))$ is, with high probability, close to the value $\mathcal{A}(\varphi(\bar{a}))$ which is determined by the values $\mathcal{A}(\varphi_{i}(\bar{a}))$, $i=1,\ldots,k$, where $\varphi_{i}\in L_{0}$ are 0/1-valued. In earlier work like [9, 14, 15] and [12, 13] (if the later are reformulated in the context of $PLA^{*}$) $L_{0}$ can be taken to be the set of complete consistent conjunctions of first-order literals. But this choice of $L_{0}$ is not always possible.
@@
[…]
@@
Suppose that $L_{0}\subseteq PLA^{*}(\sigma)$ and $L_{1}\subseteq PLA^{*}(\sigma)$ are $0/1$-valued and that the following hold:
@@
    For every $m\in\mathbb{N}^{+}$ and $\varphi_{j}(\bar{x},\bar{y})\in L_{0}$, for $j=1,\ldots,m$, there are $L_{\varphi_{j}(\bar{x},\bar{y})}\subseteq L_{1}$ for $j=1,\ldots,m$ such that if $\chi_{j}(\bar{x},\bar{y})\in L_{\varphi_{j}(\bar{x},\bar{y})}$ for $j=1,\ldots,m$, then there are $s,t\in\mathbb{N}^{+}$, $\theta_{i}(\bar{x})\in L_{0}$, $\alpha_{i,j}\in[0,1]$, for $i=1,\ldots,s$, $j=1,\ldots,m$, and $\chi^{\prime}_{i}(\bar{x})\in L_{0}$, for $i=1,\ldots,t$, such that for every $\varepsilon>0$ and $n$ there is $\mathbf{Y}^{\varepsilon}_{n}\subseteq\mathbf{W}_{n}$ such that $\lim_{n\to\infty}\mathbb{P}_{n}(\mathbf{Y}^{\varepsilon}_{n})=1$ and for every $\mathcal{A}\in\mathbf{Y}^{\varepsilon}_{n}$ the following hold:
    $(a)\ \mathcal{A}\models\forall\bar{x}\bigvee_{i=1}^{s}\theta_{i}(\bar{x})=1,$
@@
    $(d)\ \text{for all $i=1,\ldots,s$ and $j=1,\ldots,m$, if $\bar{a}\in(D_{n})^{|\bar{x}|}$, and $\mathcal{A}\models\theta_{i}(\bar{a})$,}$
@@
The $r$-tuple of sequences of $\bar{y}$-frequency parameters of $(\psi_{1},\ldots,\psi_{r})$ relative to $(\chi_{1},\ldots,\chi_{r})$ and $\theta_{i}$ is, by definition, $(\bar{p}_{1},\ldots,\bar{p}_{r})$, where $\bar{p}_{j}$ is the sequence of $\bar{y}$-frequency parameters of $\psi_{j}$ relative to $\chi_{j}$ and $\theta_{i}$. An $r$-tuple $(\psi_{1},\ldots,\psi_{r})$ is an $r$-tuple of $\bar{y}$-frequency parameters of $(\psi_{1},\ldots,\psi_{r})$ relative to $(\chi_{1},\ldots,\chi_{r})$ if, for some $i$, it is the sequence of $\bar{y}$-frequency parameters of $(\psi_{1},\ldots,\psi_{r})$ relative to $(\chi_{1},\ldots,\chi_{r})$ and $\theta_{i}$.
@@
(Applications of Theorem 5.9) (i) Suppose that Assumption 5.5 holds and $\varphi(\bar{x})\in PLA^{*}(\sigma)$. Part (1) of Assumption 5.5, Lemma 5.6, and Theorem 5.9, tell that $\varphi(\bar{x})$ can be “reduced”, step by step, to an asymptotically equivalent $L_{0}$-basic formula as long as, in a step when an aggregation function $F$ is to be “asymptically eliminated” we do not encounter parameters with respect to which $F$ is not up-continuous. The proofs of Lemma 5.6 and Theorem 5.9 indicate how the steps in such a reduction are carried out.
@@
is such that $\chi_{1},\ldots,\chi_{m}\in L_{1}$ (i.e. all $\chi_{i}$ are $\top$). As was explained in Example 4.13 every formula of the (0/1-valued) “probability logic” $L$ considered in [12] is equivalent to a $PLA^{*}(\sigma,L_{0},L_{1})$-formula (if $\sigma$ contains all relation symbols used by $L$). If Theorem 5.9 would hold without the continuity assumption, then every $PLA^{*}(\sigma,L_{0},L_{1})$-formula would be asymptotically equivalent to an $L_{0}$-basic formula. It would follow (as $L_{0}$ contains only quantifier-free formulas) that every formula of the logic $L$ considered in [12] is asymptotically equivalent to a quantifier-free first-order formula, and from this it would follow that for every $\varphi\in L$, $\lim_{n\to\infty}\mathbb{P}_{n}(\varphi)$ exists (see [12, Corollary 4.10]). But this is not the case, because as shown in [12, Proposition 3.1] there is $\varphi\in L$ such that the limit does not exist.
@@
Throughout the rest of this section we assume that $\psi(\bar{x},\bar{y})$ denotes the $L_{0}$-basic formula $\bigwedge_{i=1}^{s}\big{(}\psi_{i}(\bar{x},\bar{y})\to c_{i}\big{)}$, so $\psi_{i}\in L_{0}$ for all $i$. We also assume that that $\chi(\bar{x},\bar{y})\in\bigcap_{i=1}^{s}L_{\psi_{i}(\bar{x},\bar{y})}$. By Assumption 5.5 there are $\theta_{1}(\bar{x}),\ldots,\theta_{t}(\bar{x})\in L_{0}$, $\chi_{1}(\bar{x}),\ldots,\chi_{k}(\bar{x})\in L_{0}$, $\alpha_{i,j}\in[0,1]$ for $i=1,\ldots,t$ and $j=1,\ldots,s$, and $\mathbf{Y}^{\delta}_{n}\subseteq\mathbf{W}_{n}$ such that $\lim_{n\to\infty}\mathbb{P}_{n}(\mathbf{Y}^{\delta}_{n})=1$ and for all $\mathcal{A}\in\mathbf{Y}^{\delta}_{n}$,
@@
[…]
@@
$\bar{p}_{k}=\big{(}\mathcal{A}_{k}(\psi(\bar{a}_{k},\bar{b})):\bar{b}\in[n_{k}]^{|\bar{y}}\text{ and }\mathcal{A}\models\chi(\bar{a},\bar{b})\big{)},$

### v4

@@ 5. Asymptotic equivalence to basic formulas
A commonly considered context is when $D_{n}=\{1,\ldots,n\}$ and $\mathbf{W}_{n}$ is the set of all $\sigma$-structures with domain $D_{n}$. However we want to allow for more generality. For example one may be interested in $\sigma$-structures with domain $D_{n}$ such that all these structures have some common properties. For example, we may restrict some relation symbols in $\sigma$ to be interpreted as relations with degree [^1] A relation has degree at most $d$ if the Gaifman graph (also called the primal graph) corresponding to the relation has no vertex with degree more than $d$. at most 20 (say), or we could restrict a binary relation symbol $R\in\sigma$ to be interpreted as a tree, a partial order, an equivalence relation, perhaps with some additional properties. We can, but need not, require that, for some $\sigma^{\prime}\subset\sigma$ every structure in $\mathbf{W}_{n}$ has the same reduct to $\sigma^{\prime}$. If such restrictions are imposed on the $\sigma$-structures considered with domain $D_{n}$, then the sequence of cardinalities $|D_{n}|$, $n\in\mathbb{N}^{+}$, may not contain all (sufficiently large) positive integers. Note also that we do not require that $D_{n}\subseteq D_{n+1}$.
With this quite general set-up our goal is to isolate properties (given by Assumption 5.5) of “simpler” 0/1-valued sublogics $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ such that if $L_{0}$ and $L_{1}$ have these properties, then a formula $\varphi(\bar{x})\in PLA^{*}(\sigma)$ is asymptotically equivalent (see Definition 5.2) to an $L_{0}$-basic formula”, provided that every aggregation function in $\varphi$ is up-continuous (or ct-continuous) with respect to some parameters which are determined only by $(\mathbb{P}_{n}:n\in\mathbb{N}^{+})$ and the subformulas of $\varphi$. By an $L_{0}$-basic formula” we mean a formula of the form $\bigwedge_{i=1}^{k}(\varphi_{i}(\bar{x})\to c_{i})$ where $\varphi_{i}\in L_{0}$ and $c_{i}\in[0,1]$ for all $i$ and $\lim_{n\to\infty}\mathbb{P}_{n}\big(\forall\bar{x}\bigvee_{i=1}^{k}\varphi_{i}(\bar{x})\big)=1$. If some $\psi(\bar{x})\in PLA^{*}(\sigma)$ is asymptotically equivalent to an $L_{0}$-basic formula $\varphi(\bar{x})=\bigwedge_{i=1}^{k}(\varphi_{i}(\bar{x})\to c_{i})$, then, for sufficiently large $n$, $\mathcal{A}\in\mathbf{W}_{n}$ and $\bar{a}\in(D_{n})^{|\bar{x}|}$, the value $\mathcal{A}(\psi(\bar{a}))$ is, with high probability, close to the value $\mathcal{A}(\varphi(\bar{a}))$ which is determined by the values $\mathcal{A}(\varphi_{i}(\bar{a}))$, $i=1,\ldots,k$, where $\varphi_{i}\in L_{0}$ are 0/1-valued. In earlier work like [10, 17, 18] and [13, 14] (if the later are reformulated in the context of $PLA^{*}$) $L_{0}$ can be taken to be the set of complete consistent conjunctions of first-order literals. But this choice of $L_{0}$ is not always possible.
@@
The intuition behind the next assumption is that for some sets of formulas $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ and all $\varphi(\bar{x},\bar{y})\in L_{0}$ there is a set $L_{\varphi(\bar{x},\bar{y})}\subseteq L_{1}$ of formulas defining some “allowed” conditions, and if $\chi(\bar{x},\bar{y})\in L_{\varphi(\bar{x},\bar{y})}$, then the fraction $|\varphi(\bar{a},\mathcal{A})\cap\chi(\bar{a},\mathcal{A})|/|\chi(\bar{a},\mathcal{A})|$ is with high probability close to a number $\alpha$ that depends only on the involved formulas and the sequence $\mathbb{P}_{n}$. The assumptions in the main results of for example [10, 14, 17, 18] imply that Assumption 5.5 below holds if $L_{0}$ is the set of all formulas which are (consistent) conjunctions of first-order literals (and more generally quantifier-free first-order formulas). In the same articles $L_{1}$ can, depending on the result, be either $\{\top\}$ (where, for any $\bar{x}$ and $\bar{y}$, we can view $\top$ as a formula $\chi(\bar{x},\bar{y})$ which has value 1 for any $\bar{a}$ and $\bar{b}$ from any structure), the set of all complete and consistent conjunctions of identities ($x=y$) and nonidentities ($x\neq y$), or the set of all consistent conjuctions of identities and nonidentities; in the same results $L_{\varphi(\bar{x},\bar{y})}=L_{1}$ for every $\varphi(\bar{x},\bar{y})\in L_{0}$.
@@
Suppose that $L_{0}\subseteq PLA^{*}(\sigma)$ and $L_{1}\subseteq PLA^{*}(\sigma)$ are $0/1$-valued and that the following conditions hold:
@@
    For every $m\in\mathbb{N}^{+}$ and all $\varphi_{1}(\bar{x},\bar{y}),\ldots,\varphi_{m}(\bar{x},\bar{y})\in L_{0}$, there are $L_{\varphi_{j}(\bar{x},\bar{y})}\subseteq L_{1}$ for $j=1,\ldots,m$ such that if $\chi_{j}(\bar{x},\bar{y})\in L_{\varphi_{j}(\bar{x},\bar{y})}$ for $j=1,\ldots,m$, then there are $s,t\in\mathbb{N}^{+}$, $\theta_{i}(\bar{x})\in L_{0}$, $\alpha_{i,j}\in[0,1]$, for $i=1,\ldots,s$, $j=1,\ldots,m$, and $\chi^{\prime}_{i}(\bar{x})\in L_{0}$, for $i=1,\ldots,t$, such that for every $\varepsilon>0$ and $n$ there is $\mathbf{Y}^{\varepsilon}_{n}\subseteq\mathbf{W}_{n}$ such that $\lim_{n\to\infty}\mathbb{P}_{n}(\mathbf{Y}^{\varepsilon}_{n})=1$ and for every $\mathcal{A}\in\mathbf{Y}^{\varepsilon}_{n}$ the following conditions hold:
    $(a)\ \mathcal{A}\models\forall\bar{x}\bigvee_{i=1}^{s}\theta_{i}(\bar{x}),$
@@
    $\text{ then }(\alpha_{i,j}-\varepsilon)|\chi_{j}(\bar{a},\mathcal{A})|\ \leq\ |\varphi_{j}(\bar{a},\mathcal{A})\cap\chi_{j}(\bar{a},\mathcal{A})|\ \leq\ (\alpha_{i,j}+\varepsilon)|\chi_{j}(\bar{a},\mathcal{A})|.$
From now on we assume that Assumption 5.5 holds, that is, we have fixed two $0/1$-valued sublogics $L_{0},L_{1}\subseteq PLA^{*}(\sigma)$ such that conditions (1) and (2) of Assumption 5.5 are satisfied.
@@
The $r$-tuple of sequences of $\bar{y}$-frequency parameters of $(\psi_{1},\ldots,\psi_{r})$ relative to $(\chi_{1},\ldots,\chi_{r})$ and $\theta_{i}$ is, by definition, $(\bar{p}_{1},\ldots,\bar{p}_{r})$, where $\bar{p}_{j}$ is the sequence of $\bar{y}$-frequency parameters of $\psi_{j}$ relative to $\chi_{j}$ and $\theta_{i}$. An $r$-tuple is an $r$-tuple of $\bar{y}$-frequency parameters of $(\psi_{1},\ldots,\psi_{r})$ relative to $(\chi_{1},\ldots,\chi_{r})$ if, for some $i$, it is the sequence of $\bar{y}$-frequency parameters of $(\psi_{1},\ldots,\psi_{r})$ relative to $(\chi_{1},\ldots,\chi_{r})$ and $\theta_{i}$.
@@
Example 5.10.
As a concrete example of an application of Theorem 5.9 let us consider the following context. Let $\sigma$ be a finite relational signature and let $\mathbf{W}_{n}$ be the set of all $\sigma$-structures with domain $\{1,\ldots,n\}$. Suppose that $\mathbb{P}_{n}$ is the uniform probability distribution on $\mathbf{W}_{n}$ for all $n$. Let $L_{0}$ be the set of all conjunctions of first-order literals (over $\sigma$) and let $L_{1}=L_{0}$. In this setting, an $L_{0}$-basic formula is one of the form $\bigwedge_{i=1}^{s}(\theta_{i}(\bar{x})\to c_{i})$ where $c_{i}\in[0,1]$ and $\theta_{i}(\bar{x})\in L_{0}$ for all $i=1,\ldots,s$, and $\forall\bar{x}\bigvee_{i=1}^{s}\theta_{i}(\bar{x})$ is a valid sentence.
One can easily prove, as in [17], that every aggregation-free formula is equivalent to an $L_{0}$-basic formula so part (1) Assumption 5.5 holds. From the proofs of the classical 0-1 law (in particular the arguments in [7]) it follows that also part (2) of Assumption 5.5 holds and we can, with the notation of that assumption, let $L_{\varphi_{j}(\bar{x},\bar{y})}=L_{1}$ for all $\varphi_{j}(\bar{x},\bar{y})\in L_{0}$. It now follows from Example 4.12 and Theorem 5.9 that if $\varphi(\bar{x},\bar{y})$ is an $L_{0}$-basic formula and $\chi(\bar{x},\bar{y})\in L_{1}$, then the formulas
$\mathrm{am}\big(\varphi(\bar{x},\bar{y}):\bar{y}:\chi(\bar{x},\bar{y})\big)\ \ \text{ and }\ \ \mathrm{gm}\big(\varphi(\bar{x},\bar{y}):\bar{y}:\chi(\bar{x},\bar{y})\big)$
(where am and gm are the arithmetic, respectively gemetric, means) are asymptotically equivalent to $L_{0}$-basic formulas. If we require that the formula called $\chi(\bar{x},\bar{y})$ above is a conjuction of formulas of the form $x=y$ or $x\neq y$, then, by the results in [17, 18], we get the same conclusion for more probability distributions (than the uniform one) and for more aggregation functions (than am and gm). In [15, 16], which use Theorem 5.9, other choices of $L_{0}$ and $L_{1}$ turn out to be appropriate for the results aimed at there. [^2] The articles [15, 16] were written after this article but published before.
@@
(Applications of Theorem 5.9) (i) Suppose that Assumption 5.5 holds and $\varphi(\bar{x})\in PLA^{*}(\sigma)$. Part (1) of Assumption 5.5, Lemma 5.6, and Theorem 5.9, tell that $\varphi(\bar{x})$ can be “reduced”, step by step, to an asymptotically equivalent $L_{0}$-basic formula as long as, in a step when an aggregation function $F$ is to be “asymptotically eliminated” we do not encounter parameters with respect to which $F$ is not up-continuous. The proofs of Lemma 5.6 and Theorem 5.9 indicate how the steps in such a reduction are carried out.
@@
is such that $\chi_{1},\ldots,\chi_{m}\in L_{1}$ (i.e. all $\chi_{i}$ are $\top$). As was explained in Example 4.13 every formula of the (0/1-valued) “probability logic” $L$ considered in [13] is equivalent to a $PLA^{*}(\sigma,L_{0},L_{1})$-formula (if $\sigma$ contains all relation symbols used by $L$). If Theorem 5.9 would hold without the continuity assumption, then every $PLA^{*}(\sigma,L_{0},L_{1})$-formula would be asymptotically equivalent to an $L_{0}$-basic formula. It would follow (as $L_{0}$ contains only quantifier-free formulas) that every formula of the logic $L$ considered in [13] is asymptotically equivalent to a quantifier-free first-order formula, and from this it would follow that for every $\varphi\in L$, $\lim_{n\to\infty}\mathbb{P}_{n}(\varphi)$ exists (see [13, Corollary 4.10]). But this is not the case, because as shown in [13, Proposition 3.1] there is $\varphi\in L$ such that the limit does not exist.
@@
Throughout the rest of this section we assume that $\psi(\bar{x},\bar{y})$ denotes the $L_{0}$-basic formula $\bigwedge_{i=1}^{s}\big(\psi_{i}(\bar{x},\bar{y})\to c_{i}\big)$, so $\psi_{i}\in L_{0}$ for all $i$. We also assume that $\chi(\bar{x},\bar{y})\in\bigcap_{i=1}^{s}L_{\psi_{i}(\bar{x},\bar{y})}$. By Assumption 5.5 there are $\theta_{1}(\bar{x}),\ldots,\theta_{t}(\bar{x})\in L_{0}$, $\chi_{1}(\bar{x}),\ldots,\chi_{k}(\bar{x})\in L_{0}$, $\alpha_{i,j}\in[0,1]$ for $i=1,\ldots,t$ and $j=1,\ldots,s$, and $\mathbf{Y}^{\delta}_{n}\subseteq\mathbf{W}_{n}$ such that $\lim_{n\to\infty}\mathbb{P}_{n}(\mathbf{Y}^{\delta}_{n})=1$ and for all $\mathcal{A}\in\mathbf{Y}^{\delta}_{n}$,
@@
$\text{ then }(\alpha_{i,j}-\varepsilon)|\chi(\bar{a},\mathcal{A})|\ \leq\ |\psi_{j}(\bar{a},\mathcal{A})\cap\chi(\bar{a},\mathcal{A})|\ \leq\ (\alpha_{i,j}+\varepsilon)|\chi(\bar{a},\mathcal{A})|.$
@@
$\bar{p}_{k}=\big(\mathcal{A}_{k}(\psi(\bar{a}_{k},\bar{b})):\bar{b}\in[n_{k}]^{|\bar{y}|}\text{ and }\mathcal{A}\models\chi(\bar{a},\bar{b})\big),$
@@
6. Conclusion
We have considered the quite expressive logic $PLA^{*}$ with truth values in the unit interval and which uses aggregation functions instead of (generalized) quantifiers. We have shown that generalized quantifiers in the sense of Mostowski [21, 22] can be expressed by appropriate aggregation functions, so $PLA^{*}$ subsumes (on finite structures) first-order logic extendend by such generalized quantifiers.
We have also studied continuity properties of aggregation functions and how the continuity properties of an aggregation function $F$ influence whether a $PLA^{*}$-formula that uses $F$ is asymptotically equivalent to a simpler formula without occurences of $F$, where ‘asymptotical equivalence’ generalizes ‘almost sure equivalence’ to the context of continuous-valued logics. The same approach has been used earlier in [17, 18], but in more specific contexts. This study has identified conditions which make sense in wider contexts than that of [17, 18] and the main result (Theorem 5.9) may therefore be useful in a variety of situations, as already exemplified by [15, 16] which use the main result of this study.
Competing interests
The authors have no competing interests as defined by Springer, or other interests that might be perceived to influence the results and/or discussion reported in this paper.
Research funding
This research did not receive funding.
Data availability
This paper does not report data generation or analysis.
Author contributions
Both authors have contributed to all results of the paper.